A quadratic expression can be factored so that (x - 12)(x + 17) = 0. What are the solutions to the equation?
Answer by ReadingBoosters(3246)

(x-12)(x+17) = 0
x-12 = 0
x=12
...
x+17=0
x=-17
...
Solutions are 12 and -17
............
Delighted to help.
-Reading Boosters
Website: wwwMyHomeworkAnswers.com
Wanting for others what we want for ourselves.